Multiply Choice:: Solve 4x - 3(x + 2) = 5(4 - x)
(A) -4 1/3
(B) 4 1/3
(C) 3/13
(D) -3/13
(E) no solution
I tried to solve this but my answer didn't match one of the answers above.

Answer by yhtong(11)

ok, heres what you do. your question is
4x - 3(x + 2) = 5(4 - x)
4x-3x-6=20-5x
solve all you can solve
x-6=20-5x
bring all the unknown to the left hand side and numbers to right hand side
x+5x=20+6
6x=26
x=26/6
x=4 1/3
therefore answer is B
student of secondary pure science stream,
YH Tong